OpenStudy (anonymous):

Little help please :D Find the x and y intercept of the polynomial function f(x)-4^2(x+2)

OpenStudy (anonymous):

I believe you mean f(x) = 4^2(x+2)?

OpenStudy (anonymous):

Im not entirely sure what im to do about the -4^2. As far as I know getting the x-intercept I would have to set (x+2)=0 But I beleive I just answered my own question. After I have to set the x's=0 I should get my Y and no, -4^2(x+2)

OpenStudy (anonymous):

nvm, read it, yes you're right.

OpenStudy (anonymous):

Sorry bout that.

OpenStudy (anonymous):

just to be clear, the equation is f(x) = -4^[2(x+2)] or f(x) = -(4^2)(x+2)

OpenStudy (anonymous):

Essentially. Though from my previous comment, am I anywhere close to the solution?